σ = reaction cross section (cm^{2}) number density of protons = \frac{ρ}{m_{p}}

L = \frac{1}{(number density) (cross section)}

= \frac{m_{p}}{ρσ}

\begin{aligned}σ(ν, p^{+}) & ∼ 10^{−50} m^{2} \\ σ(γ, p^{+}) & ∼ 10^{−29} m^{2} \end{aligned}

37Cl + ν → ^{37}Ar + e^{−} 71Ga + ν → ^{71}Ge + e^{−}

\begin{aligned}ν_{e} + n & = p^{+} + e^{−} + γ \\ ν_{μ} + n & = p^{+} + μ^{−} + γ \end{aligned}